UV light of the wavelength 800 A and 700 A falls on the hydrogen atoms in their ground state and liberates electrons with KE 1.8 eV and 4 eV respectively. Calculate Planck's constant.

Grade:11

1 Answers

Ramesh V
14 years ago

KE = hc/λ - Wf       where  Wf is work function

1.8 = hc / 8*10-8  - Wf       .       ... (1)

4 = hc / 7*10-8  - Wf                  ....(2)

2.2 = hc*108 [1/7 - 1/8]

Plancks constat : h = 4.106 x 10-15 eV-sec  or  6.578 x 10-34 J-sec

---

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch